The bond market
16. Your company owns the following bonds: Bond Market Value Duration $13 milliorn $18 million $20 million 4 If general interest rates rise from 8% to 8.5%, what is the approximate change in the value of the portfolio? (Review Chapter 3.)
Answer #1

16.

The portfolio duration is :

weight of A *duration of A + Weight of B * duration of B + Weight of C *Duration of C

= 13/51*2 + 18/51 *4 + 20/51*3

=0.2549*2 + 0.3529*4 + 0.3922*3

=0.5098 + 1.4118 + 1.1766

=3.098

Therefore, the value of the portfolio will fall by -3.098*0.005 = -0.0155 i.e -1.55%.